# Towards Human-AI Deliberation: Design and Evaluation of LLM-Empowered Deliberative AI for AI-Assisted Decision-Making

### v1

@@ Abstract.
In AI-assisted decision-making, humans often passively review AI’s suggestion and decide whether to accept or reject it as a whole. In such a paradigm, humans are found to rarely trigger analytical thinking and face difficulties in communicating the nuances of conflicting opinions to the AI when disagreements occur. To tackle this challenge, we propose Human-AI Deliberation, a novel framework to promote human reflection and discussion on conflicting human-AI opinions in decision-making. Based on theories in human deliberation, this framework engages humans and AI in dimension-level opinion elicitation, deliberative discussion, and decision updates. To empower AI with deliberative capabilities, we designed Deliberative AI, which leverages large language models (LLMs) as a bridge between humans and domain-specific models to enable flexible conversational interactions and faithful information provision. An exploratory evaluation on a graduate admissions task shows that Deliberative AI outperforms conventional explainable AI (XAI) assistants in improving humans’ appropriate reliance and task performance. Based on a mixed-methods analysis of participant behavior, perception, user experience, and open-ended feedback, we draw implications for future AI-assisted decision tool design.
AI-Assisted Decision-making, Human-AI Collaboration, Appropriate Reliance
ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
With remarkable technological advancements, AI has been increasingly used to support people in making decisions in various domains, including criminal justice (Dodge et al., 2019; Dressel and Farid, 2018), admissions (Cheng et al., 2019; Zhang et al., 2023), financial investment (Green and Chen, 2019), and medical diagnosis (Cai et al., 2019b; Lee et al., 2021), among others. Concerns surrounding AI’s accuracy, safety, ethics, and accountability (Cai et al., 2019b; Lee et al., 2021; Binns et al., 2018) have led to the widespread adoption of the AI-assisted decision-making paradigm in real-world applications (Buçinca et al., 2021; Zhang et al., 2020; Wang and Yin, 2021; Bansal et al., 2021). In this paradigm, AI performs an assistive role by providing a recommendation, while human decision-makers can choose to accept or reject AI’s suggestion in their final decision (Lai et al., 2021).
Research in recent years, however, identified two challenges within the existing AI-assisted decision-making paradigm. First, a battery of empirical studies found that people rarely trigger analytical thinking (i.e., System 2 thinking (Kahneman, 2011)) when directly presented with AI’s suggestions (Buçinca et al., 2020; Rastogi et al., 2022; Bertrand et al., 2022). As a result, people frequently over-rely on the AI’s incorrect recommendations (over-reliance) or mistakenly ignore AI’s correct suggestions (under-reliance) (Buçinca et al., 2021; Ma et al., 2023; Wang and Yin, 2021). Although some solutions have been proposed, such as displaying AI explanations (Bansal et al., 2021) and forcing people to think (e.g., applying cognitive forcing functions (Buçinca et al., 2021)), the results are mixed at best (Lai et al., 2022; Poursabzi-Sangdeh et al., 2021) (Challenge 1). Second, instead of full consensus or complete divergence, human and AI decision rationales often exhibit partial alignment (Sivaraman et al., 2023; Wang and Yin, 2023). While they may concur on certain aspects, differences may persist on others (Miller, 2023). However, in current AI-assisted decision-making systems, AI consistently provides a fixed recommendation regardless of human thoughts and humans can only accept or reject AI’s recommendation as a whole (Lai et al., 2021), with limited support for resolving conflicts or engaging in a meaningful exchange of ideas with the AI system (Miller, 2023). For example, as shown in Figure 1 (a), when the human decision-maker’s prediction is inconsistent with the AI model’s recommendation and the human only partially agrees with the AI’s reasoning (e.g., explanation), existing AI-assisted decision-making interfaces do not support any communication between humans and AI regarding conflicting opinions. This limitation may impede the effective utilization of both human and AI knowledge, hindering collaborative and complementary human-AI team performance (Challenge 2).
@@
Deliberation, characterized by thoughtful and reasoned discussion, plays a pivotal role in facilitating constructive discourse and consensus-building across various contexts (Steenbergen et al., 2003; Bächtiger and Parkinson, 2019). Deliberation proves effective in facilitating various human decision-making tasks, including deliberative politics (Black et al., 2010; Habermas, 2005; Thompson, 2008), clinical diagnosis (Schaekermann et al., 2019; Inguaggiato et al., 2019; Preisz, 2019), criminal justice (Devine et al., 2001; Van Berkel et al., 2019), and more. It offers individuals an opportunity to rigorously evaluate different perspectives, including their own, which can potentially address Challenge 1 in AI-assisted decision-making. Moreover, deliberation allows participants to refine their viewpoints through informed discussions about opinion discrepancies (Peirce, 2014; Good, 1950; Gough, 2007). Such a structured process may also enable humans and AI to engage in detailed discussions, potentially mitigating Challenge 2. Despite the potential benefits of deliberation, how to design mechanisms to facilitate deliberative conversation between humans and AI and how deliberations influence AI-assisted decision-making remain to be explored.
@@
In this paper, based on guidelines and theories in human deliberation (Steenbergen et al., 2003; Bächtiger and Parkinson, 2019; Bächtiger et al., 2009) and decision-making (Alvarez-Melis et al., 2021; Weed, 2005; Baumann et al., 2014), we introduce a novel framework, Human-AI Deliberation for AI-assisted decision-making (Figure 1 (b)). Instead of presenting a fixed AI suggestion for humans to accept or reject, our framework encourages humans to externalize their thoughts, enables an interactive deliberation process between humans and AI around the conflicting points of their opinions and rationales, and fosters dynamic, fine-grained updates of humans and AI’s decisions. The key component of this framework is Deliberative AI, which has the ability to locate viewpoint dissimilarities, stimulate comprehensive deliberation with human decision-makers, and make necessary changes, even compromises, in its own suggestion as the constructive discussion unfolds. To design such an AI assistant, we propose to integrate the strength of domain-specific models (for reliable assistant information generation) and Large Language Models (LLMs, for interactivity and conversation capability). We elaborate on the design of Human-AI Deliberation framework and Deliberative AI in Section 3 and demonstrate how to instantiate the framework in an illustrative task in Section 4.
We realize that Human-AI Deliberation might be especially suitable for handling task cases that are challenging for both humans and AI. If both parties are very good at the task cases (both have high decision accuracy) or one party is very proficient in the task cases, deliberation may not be necessary. Therefore, we study the impact of Human-AI Deliberation when the task case is challenging for both parties. Specifically, using our proposed Human-AI Deliberation as a research probe, we aim to explore the following research questions.
When dealing with task cases that are challenging for both humans and AI:
@@
  • RQ1: How will Human-AI Deliberation affect task performance and humans’ reliance (and its appropriateness) on AI suggestions?
@@
  • RQ2: How will Human-AI Deliberation affect humans’ perceptions of the AI and user experience?
  • RQ3: How will humans engage in the deliberation process when working with Deliberative AI?
@@
  • RQ4: How will humans perceive the effectiveness of the proposed Human-AI Deliberation and what can be improved for future Human-AI Deliberation design?
@@
To answer these questions, we conducted an exploratory study in the context of an illustrative task (college graduate admission). We recruited participants with graduate admissions experience (at least once admitted to a graduate program) on Prolific and asked them to predict an applicant’s chance of getting an offer based on the applicant’s profile. Based on a pilot study, we selected some task cases that were edge cases and therefore challenging for both humans and AI (e.g., their accuracy on these task cases was around 50-55%). We compared the proposed Deliberative AI with traditional explainable AI (XAI) and human alone baselines. Our experimental results revealed that Human-AI Deliberation has the potential to enhance decision accuracy and promote appropriate reliance on AI recommendations compared to traditional XAI assistants. We then derived valuable insights by analyzing participants’ questionnaire ratings, human behaviors, conversation logs, and open-ended feedback.
@@
  • We propose a novel conceptual framework, Human-AI Deliberation, grounded in decision-making and deliberation theories, for humans and AI to deliberate their conflicting thoughts in a decision task. Accordingly, we design the architecture of Deliberative AI that is capable of deliberation with humans through a novel integration of domain-specific models and LLMs.
  • We demonstrate the instantiation of the framework and Deliberative AI in an illustrative task (college graduate admission), including the implementation of different components and interface design.
  • We conduct an exploratory study to gain an initial understanding of how Human-AI Deliberation might impact the decision-making process and how humans would perceive this novel AI assistance. Additionally, we demonstrate its potential to improve decision accuracy and promote appropriate human reliance on AI.
@@ 2. Related Work
2.1. AI-Assisted Decision-Making: Inadequacies and Challenges
Artificial Intelligence (AI) is increasingly used in decision-making across various domains (Ma et al., 2019, 2022b; Dastin, 2018; Dilsizian and Siegel, 2014; Khandani et al., 2010; Zhu and Ma, 2019). However, AI’s real-world applications are not infallible, still far from 100% accuracy (Shi et al., 2023; Gao et al., 2018; Ma et al., 2022a). This is especially concerning in high-stakes domains like medicine and criminal justice, leading to ethical and legal complexities (Cai et al., 2019b; Lee et al., 2021; Binns et al., 2018). To address this, the prevalent paradigm of AI-assisted decision-making has emerged, drawing substantial attention in the Human-Computer Interaction (HCI) and AI communities (Buçinca et al., 2021; Zhang et al., 2020; Wang and Yin, 2021; Bansal et al., 2021). In this paradigm, AI takes on a supportive role, offering recommendations for human decision-makers to accept or reject in their final decisions (Lai et al., 2021).
Research in AI-assisted decision-making encompasses various objectives, such as team performance, decision fairness, efficacy, efficiency, understanding of AI, trust, reliance on AI, and subjective experience (Zhang et al., 2020; Ma et al., 2023; Chiang et al., 2023; Wang et al., [n.d.]; Bansal et al., 2019a; Dressel and Farid, 2018; Wang and Yin, 2021; Cheng et al., 2019; Buçinca et al., 2021; Ma et al., 2023; Yang et al., 2020; Schemmer et al., 2023; He et al., 2023; Buçinca et al., 2020; Gajos and Mamykina, 2022; Kaur et al., 2020; Liao and Varshney, 2021; Ma et al., 2022a). One of the key challenges in AI-assisted decision-making is achieving complementary performance, where the collaborative decision outcome surpasses what either human or AI can achieve alone (Bansal et al., 2021; Lai and Tan, 2019; Zhang et al., 2020; Schemmer et al., 2023). However, empirical studies reveal difficulties in achieving this, primarily due to two issues.
@@ 2.2. The Role of Deliberation in Human Decision Making
Effective decision-making is of paramount importance across diverse domains, and deliberation emerges as a valuable process replete with manifold advantages. First, it enhances decision quality and problem-solving (Barabas, 2004; Landemore and Page, 2015). Deliberation involves comprehensive analysis and careful evaluation of options, fostering a holistic understanding of issues by considering various factors and potential outcomes (Jacobs et al., 2009). Research underscores that deliberative decisions tend to be wiser and more effective owing to their foundation in thorough information and thoughtful analysis (Landemore, 2012; Dijksterhuis et al., 2006). Second, deliberation promotes participation and collaboration (Nabatchi and Leighninger, 2015). It encourages engagement from multiple stakeholders, thereby facilitating cooperation and communication among decision-makers (Wojcieszak et al., 2010). This collaborative process aids in resolving complex issues, providing decision-makers with ample time for dialogue and negotiations, ultimately ensuring the widespread acceptance of decisions (Fishkin, 2018). Third, deliberation has the potential to mitigate decision biases and enhance fairness (Kramer et al., 1990; Hochman et al., 2015). It assists decision-makers in objectively analyzing issues, mitigating personal biases and emotional influences, and identifying flawed perceptions and knowledge (Tetlock, 2017).
Despite the significance of deliberation in decision-making, there is a dearth of research on its integration into AI-assisted decision-making processes. To address this gap, drawing upon theories and practices in deliberation (Black et al., 2010; Habermas, 2005; Thompson, 2008; Lord and Tamvaki, 2013; Steiner et al., 2005), we propose a framework called Human-AI Deliberation to facilitate human reflection and discussion on conflicting human-AI opinions. Based on this framework, we aim to move a first step towards designing a Deliberative AI and investigating its effects on decision processes and outcomes through an exploratory empirical study.
@@ 2.3. Existing Studies on Deliberation in AI-Assisted Decision Making
Deliberation encompasses two essential facets: the analytical and contemplative consideration of issues and active discussion (Gracia, 2003). These aspects have garnered some attention in previous research on AI-assisted decision-making.
@@
Concerning the stimulation of analytical thinking in humans, one approach involves interventions aimed at encouraging deeper engagement in System 2 thinking (Kahneman, 2011) by controlling when AI information is presented. For instance, the concept of ”cognitive forcing” has explored strategies to compel human decision-makers to spend more time deliberating on problems (Buçinca et al., 2021; Park et al., 2019; Rastogi et al., 2022). Methods include asking individuals to make independent predictions before receiving AI suggestions (Ma et al., 2023; Buçinca et al., 2021) or implementing a ”slow algorithm” (Park et al., 2019; Buçinca et al., 2021). These cognitive forcing techniques have demonstrated the ability to reduce human reliance on AI. Additionally, researchers have investigated how and what information should be presented. For instance, Garhos et al. (Gajos and Mamykina, 2022) found that people engaged in more analytical thinking when AI explanations were provided without concrete recommendations. Furthermore, Danry et al. (Danry et al., 2023) introduced AI-framed Questioning to enhance critical thinking and human discernment of flawed statements. While these approaches yield positive results, they come with limitations, such as potentially leading to under-reliance and not addressing differences between human and AI perspectives.
@@
Regarding discussion, some research has explored dialogues between humans and AI in cooperative games ((FAIR)† et al., 2022; Kramár et al., 2022), but these discussions have not been specifically tailored to decision-making tasks. Recent efforts have started to integrate discussions into AI-assisted decision-making, such as Zheng et al.’s (Zheng et al., 2023) inclusion of AI in group decision-making to assess student essays. However, these efforts often rely on Wizard of Oz (WoZ) setups and lack genuinely designed AI systems for human interaction. Chiang et al. (Chiang et al., 2023) examined human-AI collaborative decision-making for recidivism risk assessment, but in their work, AI only offers suggestions and does not participate in the discussion with humans. Additionally, Zhang et al. (Zhang et al., 2023) utilize AI models as boundary objects to foster deliberation among organizations, but they do not encompass deliberation between humans and AI. Perhaps the most similar work to ours is that of Slack et al. (Slack et al., 2022). They focused on AI explanation and proposed using dialogue to improve the flexibility to support arbitrary follow-up questions that users might have and enhance people’s understanding of AI. In this paper, we also use the form of dialogue to allow people to interact with the AI, and also include the part where the AI gives explanations through dialogue (such as the AI answers users’ questions). But the key difference is that we focus on deliberation design and propose Deliberative AI which can not only “explain to users” but also actively engage users in the deliberative discussions by “asking or challenging” the users, aiming to promote people’s critical thinking.
Drawing inspiration from the principles of deliberation, we introduce a novel paradigm of Human-AI Deliberation. Additionally, we design Deliberative AI which combines the conversational capabilities of large language models (LLMs) with the predictive abilities of domain-specific models (DS-Model). Deliberative AI encourages meticulous human thinking and facilitates discussions concerning disparities between human and AI viewpoints. Through an exploratory assessment, we investigate the potential impacts of deliberation on AI-assisted decision-making. Our work takes an initial step toward designing and evaluating human-AI deliberation in decision-making contexts, offering valuable insights into the integration of deliberation into AI-assisted decision-making processes.
3. Framework: Human-AI Deliberation
In this section, we first introduce the general process of decision-making with an emphasis on the step that weighs the evidence for an alternative. Then, grounded in the theoretical construct of Weight of Evidence (WoE), we propose the Human-AI Deliberation framework. Finally, based on theories and practices in human deliberation, we introduce the design considerations and structure of Deliberative AI, the core AI component of our framework.
3.1. Unpacking Decision-Making
Decision-making is a process of making a choice from a number of alternatives to achieve a desired result (Eisenfuhr, 2011). Existing literature summarizes the general decision-making process into seven steps (U Mass, [n.d.]; Lunenburg, 2010) – Step 1: Identify the problem, Step 2: Collect information, Step 3: Identify the alternatives, Step 4: Weigh the evidence, Step 5: Choose from the alternatives, Step 6: Implement action, and Step 7: Evaluate the results.

### v2

@@ Abstract.
Traditional AI-assisted decision-making systems often provide fixed recommendations that users must either accept or reject entirely, limiting meaningful interaction—especially in cases of disagreement. To address this, we introduce Human-AI Deliberation, an approach inspired by human deliberation theories that enables dimension-level opinion elicitation, iterative decision updates, and structured discussions between humans and AI. At the core of this approach is Deliberative AI, an assistant powered by large language models (LLMs) that facilitates flexible, conversational interactions and precise information exchange with domain-specific models. Through a mixed-methods user study, we found that Deliberative AI outperforms traditional explainable AI (XAI) systems by fostering appropriate human reliance and improving task performance. By analyzing participant perceptions, user experience, and open-ended feedback, we highlight key findings, discuss potential concerns, and explore the broader applicability of this approach for future AI-assisted decision-making systems.
AI-Assisted Decision-making, Human-AI Collaboration, Deliberation, Appropriate Reliance, Large Language Models
journalyear: 2025copyright: acmlicensedconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713423isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
With remarkable technological advancements, AI has been increasingly used to support people in making decisions in various domains, including criminal justice (Dodge et al., 2019; Dressel and Farid, 2018), admissions (Cheng et al., 2019; Zhang et al., 2023), financial investment (Green and Chen, 2019), and medical diagnosis (Cai et al., 2019b; Lee et al., 2021), among others. Concerns surrounding AI’s accuracy, safety, ethics, and accountability (Cai et al., 2019b; Lee et al., 2021; Binns et al., 2018) have led to the widespread adoption of the AI-assisted decision-making paradigm in real-world applications (Buçinca et al., 2021; Zhang et al., 2020; Wang and Yin, 2021; Bansal et al., 2021). In this paradigm, AI performs an assistive role by providing a recommendation, while human decision-makers can choose to accept or reject it in their final decision (Lai et al., 2021).
@@
Research in recent years, however, identified two challenges within the existing AI-assisted decision-making paradigm. [Challenge 1] First, a battery of empirical studies found that people rarely trigger analytical thinking when directly presented with AI’s suggestions (Buçinca et al., 2020; Rastogi et al., 2022; Bertrand et al., 2022). As a result, people frequently over-rely on the AI’s incorrect recommendations (over-reliance) or mistakenly ignore AI’s correct suggestions (under-reliance) (Buçinca et al., 2021; Ma et al., 2023; Wang and Yin, 2021). Although some solutions have been proposed, such as displaying AI explanations (Bansal et al., 2021) and forcing people to think more effortfully (Buçinca et al., 2021), the results are mixed at best (Lai et al., 2022; Poursabzi-Sangdeh et al., 2021). [Challenge 2] Second, instead of full consensus or complete divergence, human and AI decision rationales often exhibit partial alignment (Sivaraman et al., 2023; Wang and Yin, 2023). While they may concur on certain aspects, differences may persist on others (Miller, 2023). However, in current AI-assisted decision-making systems, AI always provides a fixed recommendation regardless of human thoughts and humans can only accept or reject AI’s recommendation as a whole (Lai et al., 2021), with limited support for resolving conflicts or engaging in a meaningful exchange of ideas with the AI system (Miller, 2023). For example, as shown in Figure 1 (a), when the human decision-maker’s prediction is inconsistent with the AI model’s recommendation and the human only partially agrees with the AI’s reasoning (e.g., explanation), existing AI-assisted decision-making interfaces do not support any communication between humans and AI regarding conflicting opinions. This limitation may impede the effective utilization of both human and AI knowledge, hindering collaborative and complementary human-AI team performance.
@@
Deliberation, characterized by thoughtful and reasoned discussion, plays a pivotal role in facilitating constructive discourse and consensus-building across various contexts (Steenbergen et al., 2003; Bächtiger and Parkinson, 2019). Deliberation proves effective in facilitating diverse human decision-making tasks, including deliberative politics (Black et al., 2010; Habermas, 2005; Thompson, 2008), clinical diagnosis (Schaekermann et al., 2019; Inguaggiato et al., 2019; Preisz, 2019), criminal justice (Devine et al., 2001; Van Berkel et al., 2019), among others. It offers individuals an opportunity to rigorously evaluate different perspectives, including their own, which can potentially address Challenge 1 in AI-assisted decision-making. Moreover, deliberation allows participants to refine their viewpoints through informed discussions about opinion discrepancies (Peirce, 2014; Good, 1950; Gough, 2007). Such a structured process may also enable humans and AI to engage in detailed discussions, potentially mitigating Challenge 2. Despite the potential benefits of deliberation, how to design mechanisms to facilitate deliberative conversation between humans and AI and how deliberations influence AI-assisted decision-making remain to be explored.
@@
In this paper, building on established guidelines for enhancing discourse quality and identifying common ground in human deliberation (Steenbergen et al., 2003; Bächtiger and Parkinson, 2019; Bächtiger et al., 2009), as well as the weight-of-evidence approach in decision-making (Alvarez-Melis et al., 2021; Weed, 2005; Baumann et al., 2014), we propose a novel solution: Human-AI Deliberation for AI-assisted decision-making (Figure 1 (b)). Instead of presenting a fixed AI suggestion for humans to accept or reject, our approach encourages humans to externalize their thoughts, enables an interactive deliberation process between humans and AI around the conflicting points of their opinions and rationales, and fosters dynamic, fine-grained updates of humans and AI’s decisions. The key component of this approach is Deliberative AI, which has the ability to locate viewpoint dissimilarities, stimulate comprehensive deliberation with human decision-makers, and make necessary changes, even compromises, in its own suggestion as the constructive discussion unfolds. To design such an AI assistant, we propose to integrate the strength of domain-specific models (for reliable assistant information generation) and Large Language Models (LLMs, for interactivity and conversation capability). We elaborate on the architecture design of Human-AI Deliberation and Deliberative AI in Section 3 and demonstrate how to instantiate the architecture in an illustrative task in Section 4.
Since the primary purpose of deliberation is to resolve conflicts between human and AI perspectives, we intentionally selected task cases with notable human-AI disagreements for our user study, drawing from insights in our pilot study. As human-AI deliberation is designed to both resolve conflicts and mitigate inappropriate human reliance on AI, we investigate its impact on task performance and human reliance on AI. Additionally, since deliberation explicitly highlights these conflicts, we are particularly interested in examining its effect on human perceptions of AI and the overall decision-making experience. Specifically, using our proposed concept of Human-AI Deliberation as a research probe, we aim to explore the following research questions.
@@
  • RQ1: How will Human-AI Deliberation affect task performance and humans’ reliance on AI suggestions?
@@
  • RQ2: How will Human-AI Deliberation affect humans’ perceptions of the AI partner and their user experience?
@@
  • RQ3: How will humans perceive the effectiveness of the proposed Human-AI Deliberation and what can be improved for future Human-AI Deliberation design?
@@
To answer these questions, we conducted an exploratory study using a graduate admissions task. We recruited participants with graduate admissions experience (at least once admitted to a graduate program) on Prolific and asked them to predict an applicant’s chance of getting an offer based on the applicant’s profile. We compared the proposed Deliberative AI with traditional explainable AI (XAI) and human alone baselines. Our experimental results revealed that Human-AI Deliberation has the potential to enhance decision accuracy and promote appropriate reliance on AI recommendations compared to traditional XAI assistants. We conclude by discussing key implications and addressing generalizability concerns based on our design and study findings.
@@
  • We propose a novel architecture, Deliberative AI, designed to enable Human-AI Deliberation—a collaborative process where humans and AI deliberate together to resolve conflicting perspectives in decision-making tasks—by seamlessly integrating domain-specific models with large language models (LLMs).
  • We demonstrate the instantiation of the Deliberative AI in an illustrative task (college graduate admission), including the implementation of different components and interface design.
  • We conduct an exploratory study to gain an initial empirical understanding of how Human-AI Deliberation might impact the decision-making process and how humans would perceive this novel AI assistance. Additionally, we demonstrate its potential to improve decision accuracy and promote appropriate human reliance on AI.
@@ 2. Related Work
2.1. AI-Assisted Decision-Making: Objectives and Challenges
Artificial Intelligence (AI) is increasingly used in decision-making across various domains (Ma et al., 2019, 2022b; Dastin, 2018; Dilsizian and Siegel, 2014). However, AI’s real-world applications are not infallible, still far from 100% accuracy (Shi et al., 2023; Gao et al., 2018; Ma et al., 2022a). This is especially concerning in high-stakes domains like medicine and criminal justice, leading to ethical and legal complexities (Cai et al., 2019b; Lee et al., 2021; Binns et al., 2018). To address this, the prevalent paradigm of AI-assisted decision-making has emerged, drawing substantial attention in the Human-Computer Interaction (HCI) and AI communities (Buçinca et al., 2021; Zhang et al., 2020; Wang and Yin, 2021; Bansal et al., 2021). In this paradigm, AI takes on an assistant role, offering recommendations for human decision-makers to accept or reject in their final decisions (Lai et al., 2021).
Research in AI-assisted decision-making spans a range of objectives, including enhancing team performance (Zhang et al., 2020; Ma et al., 2023), promoting decision fairness (Chiang et al., 2023; Wang et al., [n.d.]), improving efficacy and efficiency (Cheng et al., 2019), fostering understanding of AI (Wang and Yin, 2021; Cheng et al., 2019), building trust and appropriate reliance on AI (Schemmer et al., 2023; He et al., 2023), and enriching subjective user experiences (Liao and Varshney, 2021; Ma et al., 2022a). Among these, a key goal is achieving complementary performance—where the collaborative outcomes of human-AI teams exceed what either humans or AI could achieve independently (Bansal et al., 2021). Despite its importance, recent empirical studies highlight persistent difficulties in reaching this goal (Zhang et al., 2020; Bansal et al., 2021; Rastogi et al., 2022), driven by two primary challenges.
@@ 2.2. The Role of Deliberation in Human Decision Making
Effective decision-making is crucial across various domains, and deliberation offers significant advantages. It improves decision quality and problem-solving by enabling comprehensive analysis and evaluation of options, fostering a deeper understanding of issues and outcomes (Barabas, 2004; Landemore and Page, 2015). Deliberative decisions are often wiser and more effective due to their basis in thorough analysis (Landemore, 2012; Dijksterhuis et al., 2006). Additionally, deliberation promotes participation and collaboration, encouraging stakeholder engagement and facilitating communication, which aids in resolving complex issues and ensuring decision acceptance (Nabatchi and Leighninger, 2015; Wojcieszak et al., 2010; Fishkin, 2018). Finally, it helps mitigate decision biases and enhance fairness by enabling objective analysis and reducing emotional influences (Kramer et al., 1990; Hochman et al., 2015; Tetlock, 2017).
Despite the significance of deliberation in decision-making, there is a dearth of research on its integration into AI-assisted decision-making processes. To address this gap, drawing upon theories and practices in deliberation (Black et al., 2010; Habermas, 2005; Thompson, 2008; Lord and Tamvaki, 2013; Steiner et al., 2005), we propose Human-AI Deliberation to facilitate human reflection and discussion on conflicting human-AI opinions. Based on this approach, we aim to move a first step towards designing a Deliberative AI and investigating its effects on decision processes and outcomes through an exploratory empirical study.
@@ 2.3. Existing Studies on Deliberation in AI-Assisted Decision Making
Deliberation enhances decision-making by integrating diverse perspectives, improving solution quality, and fostering critical thinking (Gracia, 2003). It involves analytical reflection and active discussion(Gracia, 2003), both of which have been explored separately in research on AI-assisted decision making.
@@
To stimulate analytical thinking, different interventions have been designed to encourage deeper engagement with System 2 thinking (Kahneman, 2011), such as “cognitive forcing” techniques that prompt more deliberation. Examples include asking individuals to make independent predictions before receiving AI input (Buçinca et al., 2021; Park et al., 2019; Rastogi et al., 2022) or using “slow algorithms” to reduce reliance on AI. Additionally, providing AI explanations without concrete recommendations (Gajos and Mamykina, 2022) and using AI-framed questioning (Danry et al., 2023) have been shown to enhance critical thinking. However, these approaches may lead to under-reliance and may not fully address differences between human and AI perspectives.
@@
Some studies have explored human-AI dialogues in cooperative games ((FAIR)† et al., 2022; Kramár et al., 2022), but these were not tailored to decision-making tasks. Recent work has begun integrating discussions into AI-assisted decision-making. For example, Zheng et al. (Zheng et al., 2023) included AI in group decisions for student essay evaluations, though these efforts often rely on Wizard of Oz setups rather than purpose-built AI systems. Similarly, Chiang et al. (Chiang et al., 2023) studied collaboration between AI and two humans in recidivism risk assessment but limited the AI’s role to providing suggestions without active discussion. Other efforts, such as Zhang et al. (Zhang et al., 2023), used AI models as tools to facilitate deliberation among organizations but did not address direct deliberation between humans and AI. Perhaps the most relevant work is by Slack et al. (Slack et al., 2022), who explored dialogue-based AI explanations to handle follow-up user questions and improve understanding. However, the key difference is that we focus on deliberation design and propose Deliberative AI which can not only “explain to users” but also actively engage users in the deliberative discussions by “asking or challenging” the users, aiming to promote people’s critical thinking.
Previous work has touched upon the idea of having AI serve deliberation among humans, but to our knowledge, no research has directly facilitated deliberation between humans and AI. Our work takes an initial step toward designing and evaluating human-AI deliberation in decision-making contexts, offering valuable insights into integrating deliberation into AI-assisted decision-making processes.
3. Deliberative AI for Human-AI Deliberation
This section introduces deliberation into the decision-making process, focusing on weighing the evidence. Building on the Weight of Evidence (WoE) framework, we propose a Human-AI Deliberation architecture and present the design considerations and architecture of Deliberative AI, an AI assistant capable of engaging in deliberation with humans.
3.1. Integrating Deliberation into Decision-Making
Decision-making involves selecting the best choice from a range of alternatives to achieve a desired outcome (Eisenfuhr, 2011). The process is typically summarized in seven steps [^1] Seven-step decision making: Step 1: Identify the problem, Step 2: Collect information, Step 3: Identify the alternatives, Step 4: Weigh the evidence, Step 5: Choose from the alternatives, Step 6: Implement action, and Step 7: Evaluate the results., with weighing the evidence being a critical step due to its direct influence on subsequent decision outcomes (U Mass, [n.d.]; Lunenburg, 2010). We propose conducting Human-AI Deliberation during this phase because disagreements often arise between human perspectives and AI suggestions in this phase and these human-AI disagreements can lead to conflicting opinions and divergent outcomes in later steps (Lai et al., 2021).